OpenStudy (anonymous):

simplify \[\frac{ 2x ^{-3} }{ -3y ^{-2} }\]

OpenStudy (anonymous):

do you know what is x^-3/x^-2

OpenStudy (anonymous):

x^2/x^3?

OpenStudy (anonymous):

\[\frac{x ^{A}}{x^{B}} = x^{A-B}\]

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

yeah. So now can you solve the problem?

OpenStudy (anonymous):

what about the different bases?

OpenStudy (anonymous):

they're not different

OpenStudy (anonymous):

my bad, I need to amend the question.

OpenStudy (anonymous):

\[2x^{A} =/= (2x)^{A}\]

OpenStudy (anonymous):

What's there to simplify ?

OpenStudy (anonymous):

you can't really simplify it, but there's one thing you can do: switching the place on the fraction and making the exponent positive

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the rule of that?

OpenStudy (anonymous):

\[a/x = a * (1/x)\]

OpenStudy (anonymous):

and 1/(1/x) = x

OpenStudy (anonymous):

a/(b/x) = ax/b

OpenStudy (anonymous):

so if an exponent is negative, you just switch the place in the fraction and make the exponent positive

OpenStudy (anonymous):

cheers.

OpenStudy (anonymous):

np